# LongGenbench: Benchmarking Long-Form Generation in long-context LLMs

### v3

@@ Abstract
The abilities of long-context language models (LMs) are often evaluated using the ”Needle-in-a-Haystack” (NIAH) test, which comprises tasks designed to assess a model’s ability to identify specific information (”needle”) within large text sequences (”haystack”). While these benchmarks measure how well models understand long-context input sequences, they do not effectively gauge the quality of long-form text generation—a critical aspect for applications such as design proposals and creative writing. To address this gap, we have introduced a new long-form text evaluation benchmark, LongGenBench, which tests models’ ability to identify specific events within generated long text sequences. In this benchmark, we prompt long-context LMs to create long-form text that must include particular events or constraints and evaluate their ability to incorporate these elements. We evaluated ten long-context LMs across four distinct scenarios, three types of prompt instructions, and two different generation-length settings (16K and 32K). Although these models perform well on NIAH benchmarks, none demonstrated satisfactory performance on the LongGenBench , raising concerns about their ability to generate coherent long-form text that follows instructions. Additionally, as the length of the generated text increases, all models exhibit a significant drop in performance [^1] We have open-sourced LongGenBench to support the comprehensive evaluation of long-context LMs in long-form text generation. Our code and data are available at https://github.com/mozhu621/LongGenBench/..
@@ 1 Introduction
Recent advances in large language models (LLMs) have significantly improved their ability to process and generate long text sequences, which is essential for various natural language processing tasks (Xiong et al., 2023; Bai et al., 2023a). For instance, GPT-4  (Achiam et al., 2023) manages context windows of up to 128K tokens, Claude 2.1 (Anthropic, 2024a) supports up to 200K tokens, Claude 3 series (Anthropic, 2024b) handles inputs exceeding 1 million tokens, and Gemini 1.5 (Reid et al., 2024) processes context windows that span several million tokens. As LLMs become increasingly capable of processing long input text sequences, various methods have been proposed to evaluate their effectiveness in handling extended contexts. Existing benchmarks employ synthetic or hybrid evaluation, such as passkey retrieval (Mohtashami & Jaggi, 2023), needle-in-a-haystack (Kamradt, 2023), Longbench (Bai et al., 2023a), Ruler (Hsieh et al., 2024), and Needlebench (Li et al., 2024), to evaluate the retrieval, understanding, and reasoning capabilities of LLMs. However, currently there are no automated benchmarks for evaluating long output text sequences, which overlooks the quality aspect in long-form text generation tasks.
Generating long-form text sequences is crucial for various natural language processing tasks, such as document summarization(Kumar et al., 2024), creative writing(Hua & Wang, 2020; Hu et al., 2022) and comprehensive question answering(Stelmakh et al., 2022; Lee et al., 2023; Bai et al., 2024). However, evaluating these long sequences poses significant challenges. First, there is a lack of evaluation datasets, as most existing ones are designed for shorter text output ($\leq$ 2K tokens) (Fan et al., 2018; 2019a; Dasigi et al., 2021), making them less suitable for super-long-form generation  ($\geq$ 16K tokens) (Bai et al., 2024). Furthermore, as the input context window gets increasingly longer, long-form generation tasks will naturally produce even longer sequences, which amplifies the gap found in the existing dataset. Second, there is no robust method for reliably assessing the quality of these long text sequences. Researchers often rely on manual evaluations, which, while thorough, are costly and impractical for large-scale assessments. Alternatively, researchers resort to the approach “LLM-as-a-judge” (Zheng et al., 2024), which uses models capable of processing long inputs to evaluate text quality. However, the scores generated by these black-box models can be difficult to explain and may not align with human expectations, raising concerns about their validity. These challenges highlight the need for a more specialized dataset with reliable evaluation methods for super-long-form text generation.
This study introduces LongGenBench , a new benchmark assessing super-long-form generation capabilities in a long-context language model. This benchmark evaluates the quality of generated content and adherence to instructions, demanding varying levels of compliance based on the specific instructions. For instance, creative writing tasks like storytelling or diary entries for fictional characters might involve detailing significant events such as family birthdays or annual vacations. Meanwhile, project planning tasks, such as urban planning, could require the inclusion of essential constraints in the document. To establish a robust benchmark, we have designed four distinct scenarios: Diary Writing, Menu Design, Skyscraper Design, and Urban Planning. Each scenario comprises three sub-scenarios with specific task instructions—single instance, range, and periodicity. These sub-scenarios capture and reflect the complexity behind long-form generation tasks, where long sequence texts must include and adhere to various criteria. These scenarios thus serve as a foundational base for evaluating long-form generation.
Our benchmark provides a fine-grained evaluation of generated text sequences by assessing for the presence of essential events or constraints specified in the prompt instructions. Unlike the NIAH (Kamradt, 2023), which tests the fundamental capability of long-context LMs to process lengthy input sequences and identify critical information, our benchmark focuses on instruction-following in super-long-form generation tasks. Specifically, it challenges models to include specific events or constraints, akin to verifying the needle in the output text sequence. This method not only captures and reflects the overall quality of the task but also minimizes the need for human evaluation or excessive reliance on LLM-based judgments. Our study findings indicate that current long-context language models are inadequate when handling LongGenBench, even with texts shorter than 16K (see Figure 1).
Figure 1: The left side of the figure displays the average performance results of the model at Ruler (16K) (Hsieh et al., 2024), whereas the right side shows the model’s average performance on the LongGenBench  (16K).
@@
  • To the best of our knowledge, this is the first study to introduce the challenge of super-long-form generationin long-context language models. The generation, retrieval, comprehension, and reasoning within these extended contexts are crucial.
  • We introduce LongGenBench, a dataset that includes tasks for evaluating the super-long-form generation  capabilities of LLMs across multiple text token intervals (16K and 32K) and various text depth ranges.
  • We conducted comprehensive experiments on both open-source and close-source models. Our findings indicate that most models fail to complete even a fraction of the super-long-form generation tasks they are purported to manage.
@@ 2 LongGenBenchBenchmark
In this section, we provide a detailed examination of our benchmark as outlined, encompassing Task definition, Different Scenario setups, Specific task instruction, Evaluation metric and Evaluations pipeline.
@@ 2.1 Task Definition
Evaluating the quality of super-long-form generation  is both crucial and challenging, given the inherent complexity of long texts. Traditional human evaluation methods, while precise, are expensive and not scalable. Although feasible, using large language models for assessment often lacks interpretability. Thus, we focus on the ”instruction following” task in super-long-form generation . In this study, We define a task type termed Strictly Sequential Tasks, which involves the sequential completion of subtasks $\mathbf{T}=\{T_{1},T_{2},T_{3},\ldots,T_{n}\}$, where each subtask is responsible for generating a specific volume of text. For instance, an LLM might be tasked with designing a 100-floor skyscraper, specifying the content and purpose of each floor.
@@ 2.2 Four Distinct Scenario Setups
In conjunction with the design of a 100-floor skyscraper, we have devised three distinct task scenarios to supplement our predefined tasks, as illustrated in Figure 2 (1). These scenarios fall into two categories: Temporal (Diary Writing, Menu Design) and Spatial (Skyscraper Design, Urban Planning). Moreover, each scenario incorporates both short and long version to assess the effectiveness of various output lengths. Each scenario is carefully engineered to evaluate the model’s capacity to produce context-sensitive and detailed lengthy texts tailored to our specific task requirements. Table 1 offers comprehensive descriptions for each scenario, with each designed around a unique template to generate 100 different task instructions [^2] Examples of Task Instructions for each scenario are provided in Appendix B..
Table 1: Detailed Scenario Descriptions
@@ 2.3 Specific Task Instruction
  • Single Instruction (SI): Introduces specific information uniquely within the text.
@@
  • Range Instruction (RI): Incorporates specific information at designated ranges within the text.
@@
  • Periodic Instruction (PI): Periodically disperses information at predefined intervals throughout the text.
@@
For instance, in the design of a 100-floor skyscraper, the Single Instruction may specify that the 34th floor hosts an aerial gym and the 54th floor houses a law firm. The Range Instruction might designate floors 1 through 9 as a comprehensive shopping mall, whereas the Periodic Instruction could dictate that starting from the 20th floor, every 10th floor incorporates a small aerial garden. We utilize over 20 templates for each type of instruction, with the floors or locations being randomly assigned to ensure task diversity. These settings, applied via various templates, guarantee controlled coverage across all textual positions, thus facilitating a comprehensive and efficient evaluation as illustrated in Figure 2 (2). For instance, my wife’s birthday falls on October 8th, and our family trip to Miami is scheduled from December 24th to celebrate Christmas, among other events. Through this approach, we generate the main task instructions $T$ and simultaneously acquire the corresponding $check\_set$, which supports subsequent evaluations and constructs a task conducive to super-long-form generation. Subsequently, we splice the main task prompt with the specific task instructions (STI) [^3] Each of our main task instructions $T$ splice 5 single instructions, 1 range instruction task, and 1 periodic instruction task. and add the generation prompt to form the final evalution data.
@@
Main Task Completion
@@
Specific Task Instruction Completion
To assess the performance of specific subtasks under defined conditions, we employed the Llama3.1-8B instruction model (Dubey et al., 2024) for a few-shot evaluation. This method allowed us to verify compliance of each subtask with its corresponding Specific Task Instructions, effectively transforming the assessment into a binary classification task (Completed” and ”Uncompleted). Additionally, we manually annotated 500 data points to validate the consistency of the outcomes from the Llama3.1-8B-instruct model with human assessments [^4] Detailed evaluation prompts can be found in the Appendix C.. . We then calculated the completion counts for the Single Instruction (SI), Range Instruction (RI), and Periodic Instruction (PI). To assess the performance of the SPI (Specific Task Instruction) process, we computed the ratio of the total number of relevant sub-scenarios included in the answers to the number of completed instructions, denoted as STIC-1:
@@
Additionally, we calculated the total number of sub-scenarios across all Specific Task Instructions. This enumeration supported the calculation of ratios for SI, RI, PI, to the total Specific Task Instructions, resulting in the Specific Task Instruction Completion-2 metric (STIC-2):
@@
STIC-1 is primarily concerned with the completion rate of instructions that result in sub-scenarios, focusing on whether instructions are completed. In contrast, STIC-2 assesses the overall completion of the SPI task, including the presence of sub-scenarios and their completion status.
@@
Our evaluation process consists of the following steps: 1) First, we use a long-context language model (LLM) to complete the aforementioned task instruction $T$ and obtain its Answer $A$, which is then divided into sub-tasks as $A=\{A_{1},A_{2},\ldots,A_{n}\}$. 2) Next, based on the instructions that need to be evaluated in the check_set, we identify the corresponding sub-tasks in the Answer. 3) Finally, we evaluate each sub-task by $\text{eval}(A_{i},T_{i})$ to calculate the final completion score, as specified by the algorithm 1.
@@ 3.1 Experimental Setup
Models & Inference setup
We selected ten long-context LLMs, comprising eight open-source models and two closed-source model, covering a range of model sizes from 7B to 72B with Mixture of Experts (MoE) architecture and claimed context lengths ranging from 32k to 128K tokens. Detailed specifications of these models are provided in Appendix A. We evaluated all models using the vLLM (Kwon et al., 2023) serving system, which features efficient KV cache memory management. Inference was conducted with BFloat16 precision on eight NVIDIA A800 GPUs, using greedy decoding.
Task Configurations
For each scenario, we generated 800 examples for the specified lengths (16k, 32k) according to the required chat template for each model. To ensure the models produced relevant responses without generating off-topic content or refusing to answer, we prefixed each task input with a specific answer prompt and verified the presence of the target output using recall-based accuracy measures.
@@ 3.2 Main Result
Table 2: Long-form generation Performance (%) of selected models evaluated at length from 16k and 32k. The weighted average score (wAvg) is the product of CR and STBC-2, used to represent the model’s final performance at the given task length. Note that the GPT-4-32K is currently closed for use, and the longest versions that can be used are the GPT-4o and GPT-4o-mini 16K output limitation.
Completion Rate:

### v6

# LongGenbench: Benchmarking Long-Form Generation in Long Context LLMs

@@ Abstract
Current benchmarks like “Needle-in-a-Haystack” (NIAH), Ruler, and Needlebench focus on models’ ability to understand long-context input sequences but fail to capture a critical dimension: the generation of high-quality long-form text. Applications such as design proposals, technical documentation, and creative writing rely on coherent, instruction-following outputs over extended sequences—a challenge that existing benchmarks do not adequately address. To fill this gap, we introduce LongGenBench, a novel benchmark designed to rigorously evaluate large language models’ (LLMs) ability to generate long text while adhering to complex instructions. Through tasks requiring specific events or constraints within generated text, LongGenBench evaluates model performance across four distinct scenarios, three instruction types, and two generation-lengths (16K and 32K tokens). Our evaluation of ten state-of-the-art LLMs reveals that, despite strong results on Ruler, all models struggled with long text generation on LongGenBench, particularly as text length increased. This suggests that current LLMs are not yet equipped to meet the demands of real-world, long-form text generation. We open-source LongGenBench to promote comprehensive evaluation and improvement in this critical area, with code and data available at https://github.com/mozhu621/LongGenBench.
@@ 1 Introduction
Recent advances in large language models (LLMs) have dramatically enhanced their ability to process long text sequences, supporting applications that range from document summarization to creative writing. Leading models such as GPT-4 (Achiam et al., 2023), LLaMa-3.2 (Dubey et al., 2024), and Claude 2.1 (Anthropic, 2024a) manage context windows of up to 128K tokens, with the Claude 3 series (Anthropic, 2024b) handling inputs exceeding 1 million tokens. However, while much attention has been given to these models’ ability to retrieve and understand long input text sequences, far less focus has been placed on their ability to generate coherent and high-quality long-form text outputs—a critical requirement for tasks such as design proposals and creative writing.
Long-form text generation is crucial for real-world applications that require detailed, well-structured narratives, such as document summarization (Kumar et al., 2024), creative writing (Hua & Wang, 2020; Hu et al., 2022), and comprehensive question answering (Stelmakh et al., 2022; Lee et al., 2023; Bai et al., 2024). Despite this importance, current benchmarks are limited in their ability to assess long-form generation, focusing instead on shorter text outputs ($\leq$ 2K tokens) (Fan et al., 2018; 2019a; Dasigi et al., 2021), making them unsuitable for tasks requiring outputs of $\geq$16K tokens (Bai et al., 2024). The challenge is further compounded by the lack of robust methods for evaluating these long sequences. The ability to follow instructions is essential for long text generation (Reversed NIAH [^1] Analogous to NIAH, which involves searching for a needle (retrieval) within a long input, the reversed NIAH entails placing a specific needle (instruction-following) at a designated position within a long output.), just as effective information retrieval is fundamental for processing long-context inputs (NIAH(Kamradt, 2023)). However, current benchmarks do not adequately assess whether the generated text adheres to the specific directives of a prompt. For instance, a prompt may require incorporating specific information at a certain point in a lengthy document, but evaluations often fail to verify the model’s compliance with such instructions. This oversight represents a significant shortcoming in benchmarking, particularly because performance under explicit constraints typically predicts outcomes in tasks with more implicit constraints, such as story generation or academic paper production. If a model struggles with explicit requirements, it is likely to underperform in scenarios with subtler constraints.
Manual evaluations, while thorough, are both costly and impractical at scale. Meanwhile, automated evaluations using ”LLM-as-a-judge” methods (Zheng et al., 2024) often yield results that are difficult to interpret and may not align with human judgments, raising concerns about their reliability. This highlights the need for more specialized benchmarks capable of reliably assessing the quality of super-long-form text generation.
To address this gap, we present LongGenBench, a novel benchmark designed to evaluate the quality of super-long-form text generated by long-context LLMs. Unlike existing benchmarks that primarily test retrieval or reasoning over long inputs, LongGenBench  focuses on the model’s ability to generate content that follows complex instructions over extended sequences. Our benchmark introduces tasks that reflect real-world generation challenges, such as diary writing, menu planning, and urban design, where the text must adhere to specific constraints provided in the prompt. These tasks assess whether models can correctly incorporate specific details at designated points in the text, ensuring the generated content meets the requirements laid out in the prompt. By evaluating texts up to 32K tokens, LongGenBench  is the first benchmark to systematically test the ability to generate instruction-compliant long-form content across extended lengths. Table 1 summarizes the different benchmarks supporting long-context retrieval and generation tasks.
The evaluation tasks are organized into four distinct scenarios: Diary Writing, Menu Design, Skyscraper Design, and Urban Planning, each with varying complexity. The scenarios involve sub-tasks such as single instance, range, and periodicity, simulating realistic constraints that a model must account for. This setup allows us to measure the model’s ability to generate detailed, contextually rich outputs that satisfy a wide array of criteria.
@@
  • To the best of our knowledge, this is the first study to address the challenge of super-long-form generation in long-context language models, highlighting the critical importance of generating coherent, high-quality text in extended contexts.
  • We introduce LongGenBench , a comprehensive dataset that provides a diverse set of tasks specifically designed to evaluate the super-long-form generation capabilities of LLMs across varying token lengths (16K and 32K) and levels of text complexity.
  • We perform extensive experiments on both open-source and closed-source models, revealing that despite their advanced capabilities, most models struggle significantly with super-long-form generation tasks, particularly in maintaining instruction adherence and coherence over long outputs.
Table 1: Comparison of Long-Context LLM Benchmarks. For the retrieval tasks’ datasets, we measure length based on the number of processing tokens, while for the generation tasks’ datasets, we calculate the average number of generation words produced by LLMs. ‘Long-length’ indicates if LLMs to analyze or generate text that is at least 8K token.
Type of Task Benchmark Type of Data Avg Len Long-Length
Retrieval Longbench(Bai et al., 2023) hybrid $\sim$8k
NIAH(Kamradt, 2023) synthetic Any
Ruler(Hsieh et al., 2024) synthetic Any
Generation ELi5(Fan et al., 2019b) hybrid $\sim$0.2K
Longwrite(Bai et al., 2024) synthetic $\sim$2.7K
LongGenBench(Ours) synthetic $\sim$20K
@@ 2.1 Task Definition
Evaluating the quality of super-long-form generation  presents a unique set of challenges due to the inherent complexity of long texts. Traditional human evaluation methods, while precise, are expensive and not scalable. Although using large language models for assessment is feasible, their lack of interpretability often hampers their utility. Thus, we focus on the ”instruction-following” task in super-long-form generation, where the most must include specific details in the generated text. This task reflects real-world scenarios that require a high degree of attention to detail over extended sequences, such as technical documentation or detailed design proposals. In this study, we define a task type termed Strictly Sequential Tasks, which involves the sequential completion of subtasks $\mathbf{T}=\{T_{1},T_{2},T_{3},\ldots,T_{n}\}$, where each subtask is responsible for generating a specific volume of text. For instance, an LLM might be tasked with designing a 100-floor skyscraper, specifying the content and purpose of each floor.
@@ 2.2 Four Distinct Scenario Setups
To comprehensively assess the long-form generation capabilities of models, we have devised four distinct task scenarios to supplement our predefined tasks, as illustrated in Figure 1 (1). These scenarios fall into two categories: Temporal (Diary Writing, Menu Design) and Spatial (Skyscraper Design, Urban Planning). Moreover, each scenario incorporates both short and long versions to assess the effectiveness of various output lengths.
These scenarios were carefully chosen to reflect both creative and technical long-form generation tasks. They offer a diverse set of challenges by including temporal tasks (e.g., Diary Writing) that require maintaining consistent information over time and spatial tasks (e.g., Urban Planning) that test the model’s ability to handle spatial relationships and detailed designs. These scenarios mirror real-world applications, from planning documents to creative writing, and thus provide a comprehensive evaluation of long-context LLMs. Table 2 offers comprehensive descriptions for each scenario, with each designed around a unique template to generate 100 different task instructions [^2] Examples of Task Instructions for each scenario are provided in Appendix A..
Table 2: Scenario task descriptions
@@ 2.3 Specific Task Instruction
  • Single Instruction (SI): Injects specific information at a unique point within the generated text.
@@
  • Range Instruction (RI): Requires the model to incorporate information within specified ranges of the text.
@@
  • Periodic Instruction (PI): Distributes information at predefined intervals throughout the text.
@@
For example, in the design of a 100-floor skyscraper, the Single Instruction may specify that the 34th floor hosts an aerial gym and the 54th floor houses a law firm. The Range Instruction might designate floors 1 through 9 as a comprehensive shopping mall, whereas the Periodic Instruction could dictate that starting from the 20th floor, every 10th floor incorporates a small aerial garden.
We utilize over 20 templates for each type of instruction, with the floors or locations being randomly assigned to ensure task diversity. These settings, applied via various templates, guarantee controlled coverage across all textual positions, thus facilitating a comprehensive and efficient evaluation, as illustrated in Figure 1 (2).
Through this approach, we generate the main task instructions $T$ and simultaneously acquire the corresponding $Check\_set$, which supports subsequent evaluations and constructs a task conducive to super-long-form generation. Subsequently, we splice the main task prompt with the specific task instructions (STI) [^3] Each of our main task instructions $T$ splice 5 single instructions, 1 range instruction task, and 1 periodic instruction task. and add the generation prompt to form the final evaluation data.
@@
Main Task Completion.
@@
Specific Task Instruction Completion (STIC-1).
This metric evaluates the model’s adherence to specific task instructions. We calculate the completion counts for the Single Instruction (SI), Range Instruction (RI), and Periodic Instruction (PI). STIC-1 quantifies how well the model follows these instructions across subtasks, focusing on whether the instructions are correctly implemented. For example, in the Skyscraper Design task, if the model is instructed to place an aerial gym on the 34th floor and consistently places it on a different floor, it would receive a lower STIC-1 score.
@@
Specific Task Instruction Completion-2 (STIC-2).
STIC-2 provides a more granular assessment by measuring the overall completion of specific task instructions, including their presence and execution quality across all subtasks. In addition to adherence, it assesses whether the model consistently follows these instructions throughout the entire task. For instance, if the model periodically repeats certain elements but not at the required intervals, it would affect its STIC-2 score.
@@
STIC-1 is primarily concerned with the completion rate of instructions that result in sub-scenarios, focusing on whether instructions are correctly executed. In contrast, STIC-2 assesses the overall completion of the specific instruction task, including the presence of sub-scenarios and their completion status.
@@
Our evaluation process follows a structured pipeline: First, we use a long-context LLM to complete the task instruction $T$, generating an answer $A$, which is then divided into sub-tasks as $A=\{A_{1},A_{2},\ldots,A_{n}\}$. Next, based on the specific instructions in the check_set, we identify the relevant sub-tasks within $A$. Finally, we evaluate each sub-task by $\text{eval}(A_{i},T_{i})$ to compute the final completion score, as detailed in Algorithm 1. This pipeline ensures that the evaluation is both systematic and comprehensive, assessing the model’s performance across different instruction settings and levels of complexity. While LongGenBench primarily evaluates the model’s ability to follow detailed instructions, future work could expand the benchmark to include more open-ended tasks that assess creativity and logical reasoning. This would provide a broader evaluation of a model’s capabilities in generating coherent, engaging, and logically sound long-form text.
@@ 3.1 Experimental Setup
Models.
We selected ten long-context large language models (LLMs), comprising eight open-source and two closed-source models. These models range in size from 7B to 72B parameters, with one featuring a Mixture of Experts (MoE) architecture. The claimed context lengths of these models vary from 32K to 128K tokens [^4] Detailed specifications of these models are provided in Appendix C.. These models were selected to represent a diverse array of architectures, covering both Mixture of Experts and standard transformer designs, as well as a range of parameter sizes. This diversity ensures a comprehensive evaluation of their ability to handle long-context tasks.
Inference Setup.
We utilized the vLLM (Kwon et al., 2023) system, which optimizes key-value (KV) cache memory for efficient large-scale inference. This system is crucial for handling long-form generation efficiently, reducing memory overhead, and maximizing inference throughput. Inferences were performed using BFloat16 precision on 8×8$\times$8 × NVIDIA A800 GPUs, employing greedy decoding to generate the outputs. This setup ensured consistency and efficiency in the inference process.
Task Configurations.
For each scenario, we generated 800 examples at two specified lengths: 16K tokens and 32K tokens. The generation was based on designated templates for each model, ensuring task-specific relevance. The tasks were selected to reflect both creative and technical long-form generation challenges, such as diary writing, urban planning, and skyscraper design. To ensure the relevance of the generated content and prevent off-topic responses or refusals to answer, we prefixed each task input with a carefully curated answer prompt designed to guide the model’s output. The tasks were specifically selected to test the models’ ability to generate instruction-following long-form content in both creative and technical contexts. For example: In the Urban Planning task, models were tasked with generating a detailed plan for a new urban district, including descriptions of key facilities such as parks, schools, and transportation systems.
Evaluation Metric.
We evaluated model performance using the three metrics defined in Section 2.4: Main Task Completion, Specific Task Instruction Completion-1 (STIC-1), and Specific Task Instruction Completion-2 (STIC-2). These metrics provided a comprehensive assessment of the models’ ability to adhere to instructions and generate coherent long-form text.
@@ 3.2 Main Result
Table 3: Long-form generation Performance (%) of selected models evaluated at length from 16k and 32k. The weighted average score (wAvg) is the product of CR and STIC-2, used to represent the model’s final performance at the given task length. Note that the GPT-4-32K is currently closed for use, and the longest versions that can be used are the GPT-4o and GPT-4o-mini 16K output limitation.
@@
Results.
The results of the long-form text generation tasks for both 16K and 32K tokens are summarized in Table 3. The models exhibited significant performance variations, highlighting differences in architecture, parameter size, and training data.